Can someone please help me with the following two questions?
1. Mary just deposited $33,000 in an account paying 10% interest. She plans to leave the money in this account for seven years. How much will she have in the account at the end of the seventh year?
I got 64,307.66. Is this correct?
2. Mary and Joe would like to save up $10,000 by the end of three years from now to buy new furniture for their home. They currently have $2500 in a savings account set aside for the furniture. They would like to make equal year end deposits to this savings account to pay for the furniture when they purchase it three years from now. Assuming that this account pays 8% interest, how much should the year end payments be?
I cannot find a formula to solve this question. Can someone please help me? Greatly appreciated.

Answer by stanbon(75887)

1. Mary just deposited $33,000 in an account paying 10% interest. She plans to leave the money in this account for seven years. How much will she have in the account at the end of the seventh year?
I got 64,307.66. Is this correct?
-----------------
Since did not specify compounding I guess the money is invested at 10%
simple interest!? Is that right?
A = P(1+rt) = 33000(1+0.01*7) = 33000*(1+0.21) = $39930.00
------------------------------------------
2. Mary and Joe would like to save up $10,000 by the end of three years from now to buy new furniture for their home. They currently have $2500 in a savings account set aside for the furniture. They would like to make equal year end deposits to this savings account to pay for the furniture when they purchase it three years from now. Assuming that this account pays 8% interest, how much should the year end payments be?
I cannot find a formula to solve this question. Can someone please help me?
-------------
Formula: Bn=A(1+i)^n - (P/i)[(1+i)^n - 1]
Bn is the desired balance = $10000
A is the current amount = 2500
i is the interest rate = 0.08
n is # of years = 3
---------
You need to solve for P = periodic payment
----------
Cheers,
stan H.